A small firm builds galvanized swing sets. The investment in plant and equipment is $200,000. The variable cost per swing set is $500. The selling price of the swing set is S1000. How many swing sets would have to be sold for the firm to break even?
Answer #1
Let the firm sell x swing sets to break even.
Revenue = cost at break even point.
cost = fixed cost + variable cost
1000*x = 200000 + 500*x
500*x = 200000
x = 400
The firm will have to sell 400 units to break even.